OpenStudy (anonymous):

What is the similarity ratio of two cubes with volumes of 125 cubic meters and 216 cubic meters? Answer 125:216 5:6 25:36 1:3

OpenStudy (anonymous):

i think it is 5:6 but im not sure?

OpenStudy (ash2326):

You're right, How did you get that?

OpenStudy (anonymous):

yes you are right, volume goes with the cube

OpenStudy (anonymous):

similarity is alwasys seen in terms of one dimension... for eg. lenght breadth or radius or diameter.. and the volume is 3 dimensional so we find the ratio of the radius and then compare....

OpenStudy (anonymous):

ok thank-you!! i forgot how i worked this out it was a few lessons back but i never turned it in so im just checking all my answers before i submit it

OpenStudy (phi):

By similarity ratio they must mean the length of the side of cube 1 compared to the length of the side of cube 3. They give you the volume of each cube. You know volume = s^3 (side cubed) so find the number so that s*s*s= 125 (s=5) and s*s*s= 216 (s=6) now you take the ratio of their sides to get 5/6